# Signal Processing and Machine Learning Algorithms for Precise Timing with PICOSEC Micromegas Detectors

### v1

@@ 5.2 Validation of the ANN’s Unbiased Performance
The second training procedure approach uses as a training set the simulated pulses. Aiming for a training session unbiased of the e-peak size, the model was developed with uniform charge and amplitude distributions555The uniform region of distribution represents the vast majority of the EXP-set signals, as shown in Fig.19. Again, the ANN reaches the same timing accuracy as the full signal processing analysis of 18.5 0.6 ps.
@@ 6 Conclusions
Another technique proposed in this study is an Artificial Neural Network (ANN) for real-time signal processing, which holds great promise for fast event selection. To train the ANN, a model was created due to the need for a large training sample. After verifying that the model was consistent with the detector’s response to many photoelectrons, it was used to train the ANN to extract timing information. The ANN passed all tests with high precision timing performance, proving that it is a reliable and unbiased tool, despite the model used as the training set consisting of waveforms that are not completely uncorrelated. The emulated data proved sufficient for the ANN to learn the signal analysis procedure for precise timing. Significantly, this first implementation was developed in a regime where the detector design ensured minimal time-walk effects—only a few picoseconds in the thin-gap prototype [6]. Under such conditions, the ANN approach is highly effective. However, in detectors or conditions where time-walk contributions are more significant, the same method may no longer provide an accurate approximation, potentially limiting its timing resolution.
It should also be emphasized that the ANN technique presented here represents only the first architecture tested within this framework. Other ANN-based approaches have already been developed [19, 20] and will be reported in forthcoming publications. These complementary efforts highlight the ongoing progress in applying machine learning to PICOSEC timing, suggesting that further improvements in architecture and training strategies may unlock even higher performance.

### v3

@@ 5.2 Validation of the ANN’s Unbiased Performance
The second training procedure approach uses as a training set the simulated pulses. Aiming for a training session unbiased of the e-peak size, the model was developed with uniform charge and amplitude distributions777The uniform region of distribution represents the vast majority of the EXP-set signals, as shown in Fig.19. Again, the ANN reaches the same timing accuracy as the full signal processing analysis of 18.5 0.6 ps.
@@ 6 Conclusions
Another technique proposed in this study is an Artificial Neural Network (ANN) for real-time signal processing, which holds great promise for fast event selection. To train the ANN, a model was created due to the need for a large training sample. After verifying that the model was consistent with the detector’s response to many photoelectrons, it was used to train the ANN to extract timing information. The ANN passed all tests with high precision timing performance, proving that it is a reliable and unbiased tool, despite the model used as the training set consisting of waveforms that are not completely uncorrelated. The emulated data proved sufficient for the ANN to learn the signal analysis procedure for precise timing. Significantly, this first implementation was developed in a regime where the detector design ensured minimal time-walk effects—only a few picoseconds in the thin-gap prototype [utrobicic2024singlechannelpicosecmicromegas]. Under such conditions, the ANN approach is highly effective because the uniformity of the waveform shapes allows the network to focus on subtle temporal features rather than compensating for amplitude-dependent variations. In contrast, for detectors or operating conditions where time-walk effects are more pronounced, additional corrections or training strategies may be required, as the same method might no longer provide an accurate approximation, potentially limiting its timing resolution.
It should also be emphasized that the ANN technique presented here represents only the first architecture tested within this framework. Other ANN-based approaches have already been developed [Tsiamis:359504, charzianagnostou:343732] and will be reported in forthcoming publications. These complementary efforts highlight the ongoing progress in applying machine learning to PICOSEC timing, suggesting that further improvements in architecture and training strategies may unlock even higher performance.